1. From the give table calculate Elasticity of Price, Total Revenue and Marginal Revenue. Also, explain the relationship between AR and MR? And give conclusion too.


Price Quantity TR MR

6 0

5 100

4 200

3 300

2 400

1 500

0 600

1
Expert's answer
2021-05-18T12:37:42-0400

Solution:

Total revenue is calculated as follows:

Total Revenue (TR) = P x Q

Marginal Revenue (MR) = Divide the change in TR by the change in quantity sold.

The elasticity of price = Percentage change in the quantity demanded of a good or service divided by the percentage change in price.

 

The elasticity of price (PED) = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


Percentage change in the quantity demanded "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \times 100"


"= \frac{100 -500}{(100+500)/2 } \times 100 = \frac{-400}{300} \times 100 = -133.33\%"

 

Percentage change in the price = "\frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 } \times 100"


"=\frac{5 -1}{(5+1)/2 } \times 100 = \frac{4}{3} \times 100 = 133.33\%"

 

Price elasticity of demand (PED) = "\frac{-133.33\%}{133.33\%} = -1.0"

PED = 1

PED is, therefore, unit elastic.

 

The calculations for TR and MR are depicted in the below table:

Also, the calculation for AR is included to explain its relationship with MR.















The relationship between average revenue (AR) and marginal revenue (MR) is that when AR falls, the MR curve lies below it. This means that MR declines at a more rapid rate than AR so that the gap between AR and MR becomes wider with output increase. Although, in a perfect competition when AR is constant, MR is equal to AR. MR can both be negative or positive but AR can only be positive.

From the table above, you can see that the marginal revenue is lower than the average revenue.